OpenStudy (anonymous):

find vertex of : f(x)=2x^2+8x+12

OpenStudy (tkhunny):

Shall we "Complete the Square" or just the formula, -b/2a?

OpenStudy (anonymous):

complete the square

OpenStudy (tkhunny):

This is how I do it. I start by making it look like it needs to look when I'm done. f(x) = 2(x^2 + 4x + ______) + 12 - 2(______) The two blanks are the same. With what number shall we fill them?

OpenStudy (anonymous):

2? don't we take half of the middle term?

OpenStudy (tkhunny):

2? No. Half the middle coefficient and then square it. 4/2 = 2 2^2 = 4 Put 4 in the two blanks and simplify.